OpenStudy (anonymous):

Mr narom spend 3/5 of his money on a tv set. It cost him $432. How much money did he have left?

OpenStudy (anonymous):

Let he has initially his money as x.. So, 3/5 of x = 432 = 3x/5 = 432 So, x = 432*5/3 = 720.. So, Money left = 720 - 432 = 288.. So, he lefts with $288....

OpenStudy (saifoo.khan):

\[432 - (\frac 35 \times 432)\]

Parth (parthkohli):

omg I deleted it *WHY!?!?!*

OpenStudy (lgbasallote):

ohhh contradicting formulas >:)))

OpenStudy (anonymous):

haha it was wrong (mine)

OpenStudy (lgbasallote):

shall i be the clincher? :)

OpenStudy (lgbasallote):

all are right (even parth) @javawarrior is the only one wrong :p mwahahaha >:)) lol jk..i just dont know what you did ;)

OpenStudy (anonymous):

what are you talking about, I didnt even post a formula :) cough cough

OpenStudy (lgbasallote):

i was referring to your statement "it was wrong (mine)"

OpenStudy (lgbasallote):

im saying that that's wrong ;)

OpenStudy (anonymous):

haha @lgbasallote you clearly spend too much time on here :P